# Computational Design and Fabrication of Modular Robots with Untethered Control

### v1

@@ 3 Physical Robots
Next, we leverage the universal building block and the untethered control system to build locomotive robots. To showcase this, we built Xbot (see Fig. 4 and supplementary video: S4) with a skeletal graph consisting of 5 nodes and 4 bones, and equipped with two muscles per bone. To walk in a given direction, we actuate its hind legs (with respect to the direction) with a periodic wave similar to the motion of a caterpillar [45] using the locomotion control algorithm described in Sec. 2.
Figure 6: We demonstrate the optimization of a given frog-bot design (a) to make specialized frog-bots for their respective terrains by learning both their skeleton and gait patterns (see: Sec. 5 Exploring the design and control of terrain navigating robots). The optimized designs were then used to fabricate the frog-bots and their gaits were fine-tuned for the real-world tasks such as walking (b), climbing an inclined plane (c) and swimming (d).
@@
Finally, to showcase the vast design and control space offered by our universal building block and untethered control system, we introduce multi-terrain navigating frog-bots that can walk on a flat surface, climb an inclined plane, and swim in water. We start with a skeletal graph of 10 nodes and 12 bones and fabricate a standard frog-bot that can navigate the three terrains. We then use our design tool (see Sec. 5: Exploring the design and control of terrain navigating robots) to explore the skeletal design and control gaits and create three optimized frog-bots , one for each terrain. The fabricated frog-bots use rigid PLA material of 1mm1\ $m$ cross-sectional radius and the locomotion control algorithm described in Sec. 2 is fine-tuned for their specific terrains (see Fig. 6 and supplementary video: S6).
4 Discussion
@@
Our current framework lies in an experimental laboratory setting, which constraints the structural tessellation of the fabricated robots and restricts the range of their movement. We believe that advances in material science for 3D printing bones [46] and the synthesis of LCE-based muscles [47] along with developments in high-intensity laser actuation would help streamline the fabrication of a robot with hundreds of actuators. Engineering support that could allow us to harness the infrared radiation of the sun could enable us to control robots outside of the experimental setting [48]. Our robots provide a platform on which the community of machine learning and computational design researchers can expand upon by providing solutions for enhanced simulation that reduce the sim-to-real gap and empower them with learning-based methods [49, 50] to design robots for the future. We envision that our system will help the community make a fresh leap into a new paradigm of computationally designed modular soft robots, optimized for task-specific specifications by drawing inspiration from our proposed building block with active and untethered control.
5 Methods
@@
We 3D print the skeleton of our robots by assembling individual bones. Each bone is parameterized with length as a degree of freedom. The material used for printing, as well as the cross-sectional thickness, determine the flexibility of the material (see Appendix A 3D printing the skeleton). We characterize these bones using the ZwickRoell tensile test machine. We perform a compression test to measure the force needed to bend each bone. The loading threshold is set either to the maximum contraction length achievable by LCEs (33%) or to a force limit of 2N2\ N (twice the force LCE muscles can apply). Fig. 3 (a,b) shows the force-contraction curves for bones made of two 3D-printable materials, PLA and TPU. For each material, we analyze structures with a middle connector and without it, and three values of rod thickness. We use 1.0mm1.0\ $m$ cross section for the PLA skeleton and 2.0mm2.0\ $m$ cross section thickness for the TPU skeleton for the printed robots, as they allow preferable stiffness-to-flexibility behavior. Bones made with PLA only exhibit preferential bending due to its stiff material response, while bones fabricated with TPU showcase both preferential bending and shrinkage (see supplementary video S1)
@@
LCEs exhibit reversible deformation when heated, as a result of the rearrangement of liquid crystal molecules [42]. We fabricate rod-shaped LCEs based on the synthesis reaction reported in [43]. We enable the actuation of the muscles with electromagnetic radiations by incorporating an infrared-absorbing dye (IR820) within the polymer mixture. We conducted an infrared spectrometry test to measure the IR radiation absorption of the LCE muscles. Fig. 3 (c) shows an absorption peak around $\lambda$ = 800nm800\ nm, matching the operational wavelength of the laser and enabling the efficient transfer of heat to the LCEs. The material rods achieve a maximum contraction of 33% along the alignment direction when heated from room temperature to 120°C. We provide the synthesis protocol for the LCE muscles, their dimensions, fabrication challenges and limitations in the supplementary material (see Appendix A Synthesizing LCEs).
@@
Once the skeleton (3D-printed) and the muscles (synthesized LCEs) are ready, we attach the muscles to the skeleton using Kraftprotz super glue. An ArUco marker $8\times 8\ mm^{2}$ is placed at the joints where two bones meet. This helps in uniquely identifying the joint (see Appendix A Tracking Aruco Markers). The LCE muscles are cut with 5% extra length as slack and then glued to the endpoints at their designated location. A bottom muscle is always on the left side of the edge with the direction guided from the lower to the higher vertex index (sometimes, we manually flip the bottom and top LCE on the bone to avoid self-occlusion of bottom muscles and the information is then hard-coded for that robot). This gives us an exact geometric location of the muscles based on the position and normal values of the markers on both sides of the edge. This process is repeated for all the edges and the final robot is then ready to be used with the control system (see figure 1 (d-e-f) : skeletal graph to CAD model to fabricated robot).
@@
To build our own laser heating setup, we used a laser diode system (808nm808\ nm, 50W50\ W) manufactured by Civil Laser. The laser beam is focused using a collimator and directed to a Galvoscanner (JD2808, Sino Galvo) through a microcontroller (Teensy 4), which controls the scanner at high speeds using the XY2100XY2-100 protocol. The microcontroller processes target position data to control both the scanner and laser power. The detailed calibration process for the heating system is explained in Appendix A: Calibrating the laser heating system. Our laser control system with all its components is showcased in Fig. 7.
We implement a pulse-width-modulation algorithm on the Teensy-4 microcontroller to interpolate the intensity value between [01][0-1] and use the highest operational current of 5A5\ A with the lowest latency delay 50ms50\ ms of packet communication. Since most of the control actions rely on shrinking the muscles to a target length, we chose the laser parameters that allow us to actuate the maximum number of muscles concurrently. The ablation experiment is explained in Appendix A: Constraints on the Number of Actuators in a Robot. With this configuration of the setup, we can contract the total muscle length of up to 300mm300\ $m$. For a standard muscle of length 30mm30\ $m$, this is equivalent to approximately 30 muscles, each of which can be shrunk by maximum of 33%. This total shrink length and contraction factor define the fabrication limitations of the skeletal designs imposed by the control system (see Eq. 2).
For this configuration, we further characterize the muscle shrinkage obtained by repeating the actuation on an LCE muscle of 30mm30\ $m$. Fig. 3 (d) showcases the shrinkage behavior of muscles when activated with the laser setup for maximum intensity while varying the number of repetitions.
@@
1:Input: 3D Manifold meshes(1,n) $\Omega^{1}_{0}(V^{1}_{0},F^{1}_{0})$, $\Omega^{2}_{0}(V^{2}_{0},F^{2}_{0})$$\Omega^{n}_{0}(V^{n}_{0},F^{n}_{0})$]
2:Output: $3D$ Approx. meshes(1,n) $\Omega^{1\star}(V^{1\star},T^{\star}),\Omega^{2\star}(V^{2\star},T^{\star})...\Omega^{n\star}(V^{n\star},T^{\star})$] and 2D mesh ($\Gamma^{\star}(V^{flat\star},T^{\star})$) all having the same triangulation $T^{\star}$.
3:Energy function for the current meshes $\Omega^{i}$, $\Gamma$ and $T$.
$\mathcal{E}=\frac{1}{n}.\sum_{i=1}^{n}\Big{[}$ $\quad w_{bij}.E_{bij}(\Omega^{i})\ +w_{map}.E_{map}(\Omega^{i},\Gamma)\ +$
$\quad w_{mesh}.E_{mesh}(\Omega^{i})\ +w_{approx}.E_{approx}(\Omega^{i}_{0},\Omega^{i})\ +$
$\quad w_{boundary}.E_{boundary}(\Omega^{i})\Big{]}\ +w_{fab}.E_{fab}(\Gamma)\quad$ (1)
4:Constraints: $\forall i\in[1,n]$ and $\forall ej\in T$, and $\mathcal{L}$ be the length of edge $e_{j}$:
@@
$\forall_{e_{j}\in T}:\mathcal{L}\left(\Gamma(e_{j})\right)>30$ (2)
5:Compute the embeddings of input meshes using [51], restrict them to unit circular boundary, initialize $\Gamma^{0}$ as embedding of $\Omega^{0}_{0}$ and use $T_{0}$ as the triangulation of $\Gamma^{0}$.
6:Optimize the energy function 3 with ”Coarse” parameters from Table 2 with the constraint in Eq. 2 following Algorithm 1 as reported in [52].
7:Optimize the energy function 3 with ”Fabrication” parameters from Table 2 with the constraint in Eq. 2 following Algorithm 1 as reported in [52].
@@
The design tool aims to generate the skeletal graph for robots that can morph between different target deformations. It takes the desired target shapes as input and generates a flat skeletal graph, which after fabrication, can be actuated by the laser heating system to closely approximate the input target shapes. To obtain the skeletal graph, we approximate the target shapes with the same connectivity and ensuring that their edges are related by a scale factor. This allows our edge-shrinkage-based method to morph between them. We adapt the formulation proposed by [52] aimed at creating Surface Maps via Adaptive Triangulation for various objects, to obtain a compatible triangulation. This triangulation when embedded in 2D yields a planar skeletal graph structure. We summarize the method in Algorithm 1 The detailed modifications made to the initial formulation by [52] and the optimization steps are explained in Appendix A: Algorithm Details: Computational Design for Shape Morphing Structures. We show the efficacy of our tool by building the morph-bot (see Fig. 5), which morphs into target saddle and dome shapes with good visual accuracy. We further showcase futuristic applications that the computational design tool unlocks by fabricating finely tessellated flat-fabricated surfaces that morph into multiple target configurations (see Fig. 8).
@@
The numerical simulation uses the robot’s skeletal graph as input. We model the bones of the skeletal graph using discrete elastic rods [53]. The LCE muscles are modeled as spring actuators in the simulation, with their rest length being a function of temperature. We adjusted the material parameters of the simulation model to ensure that the motion of the base building block matches its behavior in the real world. This is done by tuning the simulation’s material parameters as described in [54]. We approximated the laser actuation control as a time-varying change of the LCEs’ rest length to its target contracted length, as given by the shape-morphing algorithm and used quasi-static simulation with contact and gravity to compute the rest-state of the robots after morphing. This allows us to verify the efficacy of the shape-morphing robots before they are manufactured. Fig. 5 (c) shows the simulation results on morph-bot .
@@
We introduce a design tool to fine-tune a given skeletal robot to achieve a desired functional task. The design space includes the length of active actuators and their control policy. We jointly optimize these two spaces to obtain a skeleton with a control scheme that performs the target function. If multiple functions are desired, we output one optimized skeleton and a unique control scheme for each function. We model the robots as rigid articulated robots and simulated them using Mujoco [55]. We simulate three different target behaviors: walking on flat terrain, climbing on an inclined plane, and swimming in water. Since Mujoco allows for fast simulation of rigid robots, we can optimize the nodal positions of the skeleton and control schemes for different target behavior using evolutionary search based algorithms. The implementation details of our optimization framework are described in Appendix A: Algorithm Details: Exploring the Design and Control of Terrain Navigating Robots. The optimized frogs exhibit the highest fitness scores for their respective target behavior. We use this tool for the designs of frog-bots . Modeling the robots as rigid robots introduces a significant sim-to-real gap, leading to discrepancy between simulated and fabricated motions. Despite this, the design tool successfully optimizes the skeletal graph of frog-bots for their specific terrains, outperforming the non-optimized frog-bot . Fig. 6 and supplementary video: S6, showcase the frog-bots designed using this tool.
6 Author information
Authors and Affiliations
Institute of Science and Technology Austria, Am Campus 1, Klosterneuburg, 3400, Austria
Manas Bhargava, Malina Strugaru, Michal Piovarci and Bernd Bickel
University of Tsukuba, 1-2 Kasuga, Tsukuba, Ibaraki 305-8550, Japan
Takefumi Hiraki
Cluster Metaverse Lab, 8-9-5 Nishigotanda, Shinagawa, Tokyo 141-0031, Japan
Takefumi Hiraki
ETH Zurich, Computational Design Laboratory, HIT H 41.3, Wolfgang-Pauli-Str. 27, Zurich, 8093, Switzerland
Michal Piovarci and Bernd Bickel
California Institute of Technology, 1200 E California Blvd, Pasadena, CA, 91125, USA
Chiara Daraio
The University of Osaka, 1-3 Machikaneyama, Toyonaka, Osaka 560-8531, Japan
Daisuke Iwai
Contributions
M.B. conducted the primary research for this study, including the synthesis of LCE material, the characterization of the 3D-printed bone structures, and the characterization of the laser heating setup with LCE muscles. M.B. also designed, fabricated, and executed experiments for the various robotic systems presented; automated the CAD design pipeline; and developed the design and simulation tools described in the paper. M.B. was responsible for drafting the original manuscript and preparing the figures. Additionally, M.B. contributed to the building and calibration of the laser setup. T.H. designed, built, and calibrated the laser actuation hardware and software and implemented the marker tracking system. T.H. also assisted in the design of the laser-based experiments. M.S. advanced the research on LCE muscles by developing the IR-absorbing LCE materials used in the study, and conducted experiments for the mechanical characterization of the bones and muscles. M.B and M.S. designed the building block of the robots and the experimental environments for the locomotive robots. M.P. contributed to the conceptualization of the design and simulation tools. B.B. conceived the study. C.D. served in an advisory role and helped with conceptualization. D.I. and B.B. supervised the project. All authors contributed to project discussions and the editing of the manuscript.

### v2

@@ Results
Next, we leverage the universal building block and the untethered control system to build locomotive robots. To showcase this, we built Xbot (see Figure 4 and supplementary video: S4) with a skeletal graph consisting of 5 nodes and 4 bones, and equipped with two muscles per bone. To walk in a given direction, we actuate its hind legs (with respect to the direction) with a periodic wave similar to the motion of a caterpillar  (?) using the locomotion control algorithm.
Figure 6: Starting with an input skeletal designs, we demonstrate the performance of our method by optimizing a frog-bot design for three different tasks: walking (A), climbing a 20-degree inclined plane (B) and swimming (C). frog-bots were optimized for their respective terrains by learning both their skeleton and gait patterns. The optimized designs were then used to fabricate the frog-bots and their gaits were hand-tuned for the real-world performance. For each terrain, we first showcase the optimized robots in the simulation. We then showcase their optimized skeleton from which the final fabricated model is obtained. Finally, we showcase the real world performance by taking their start and end snapshots.
@@
Finally, to showcase the vast design and control space offered by our universal building block and untethered control system, we introduce multi-terrain navigating frog-bots that can walk on a flat surface, climb an inclined plane, and swim in water. We start with a skeletal graph of 10 nodes and 12 bones and fabricate a standard frog-bot that can navigate the three terrains. We then use our design tool (see Exploring the Design and Control of Terrain Navigating Robots: Materials and Methods: Exploring the design and control of terrain navigating robots) to explore the skeletal design and control gaits and create three optimized frog-bots , one for each terrain. The fabricated frog-bots use rigid PLA material of 1mm1\ $m$ cross-sectional radius and the locomotion control algorithm is fine-tuned for their specific terrains (see Figure 6 and supplementary video: S6).
Discussion
@@
Our current framework lies in an experimental laboratory setting, which constraints the structural tessellation of the fabricated robots and restricts the range of their movement. We believe that advances in material science for 3D printing bones  (?) and the synthesis of LCE-based muscles  (?) along with developments in high-intensity laser actuation would help streamline the fabrication of a robot with hundreds of actuators. Engineering support that harnesses the sun’s infrared radiation could enable control of robots beyond the experimental setting  (?). Our robots provide a platform for the machine learning and computational design community to expand upon by providing enhanced simulation tools that reduce the sim-to-real gap and equip them with learning-based methods  (?, ?) to design the next generation of robots. We envision that our system will help the community make a fresh leap into a new paradigm of computationally designed modular soft robots, optimized for task-specific specifications by drawing inspiration from our proposed building block with active and untethered control.
Materials and Methods
@@
We 3D print the skeleton of our robots by assembling individual bones. Each bone is parameterized with length as a degree of freedom. The material used for printing, as well as the cross-sectional thickness, determine the flexibility of the bone (see Supplementary Information: 3D printing the skeleton). We characterize these bones using the ZwickRoell tensile test machine. We perform a compression test to measure the force needed to bend each bone. The loading threshold is set either to the maximum contraction length achievable by LCEs (33%) or to a force limit of 2N2\ N (twice the force LCE muscles can apply). Figure 3 (A,B) shows the force-contraction curves for bones made of two 3D-printable materials, PLA and TPU. For each material, we analyze structures with a middle connector and without it, and three values of rod thickness. We use 1.0mm1.0\ $m$ cross section for the PLA skeleton and 2.0mm2.0\ $m$ cross section thickness for the TPU skeleton for the printed robots, as they allow preferable stiffness-to-flexibility behavior. Bones made with PLA exhibit only preferential bending due to its stiff material response, while bones fabricated with TPU showcase both preferential bending and shrinkage (see supplementary video S1)
@@
LCEs exhibit reversible deformation when heated, as a result of the rearrangement of liquid crystal molecules  (?). We fabricate rod-shaped LCEs based on the synthesis reaction reported in  (?). We enable the actuation of the muscles with electromagnetic radiations by incorporating an infrared-absorbing dye (IR820) within the polymer mixture. We conducted an infrared spectrometry test to measure the IR radiation absorption of the LCE muscles. Figure 3 (C) shows an absorption peak around $\lambda$ = 800nm800\ nm, matching the operational wavelength of the laser and enabling the efficient transfer of heat to the LCEs. The material rods achieve a maximum contraction of 33% along the alignment direction when heated from room temperature to 120°C. We provide the synthesis protocol for the LCE muscles, their dimensions, fabrication challenges and limitations in the supplementary material.
@@
Once the skeleton (3D-printed) and the muscles (synthesized LCEs) are ready, we attach the muscles to the skeleton using Kraftprotz super glue. An $8\times 8\ mm^{2}$ ArUco marker is placed at each joint where two bones meet. This helps in uniquely identifying the joint. The LCE muscles are cut with 5% extra length as slack and then glued to the endpoints at their designated location. For each edge, a bottom muscle is always on the left side of the edge and the top muscle on the right side, with the direction guided from the lower to the higher vertex index. This gives us an exact geometric location of the muscles based on the position and normal values of the markers on both sides of the edge. This process is repeated for all the edges and the final robot is then ready to be used with the control system (see Figure 1 (d-e-f) : skeletal graph to CAD model to fabricated robot).
@@
To build our own laser heating setup, we used a laser diode system (808nm808\ nm, 50W50\ W) manufactured by Civil Laser. The laser beam is focused using a collimator and directed to a Galvoscanner (JD2808, Sino Galvo) through a microcontroller (Teensy 4), which controls the scanner at high speeds using the XY2100XY2-100 protocol. The microcontroller processes target position data to control both the scanner and laser power. The detailed calibration process for the heating system is explained in Calibrating the Laser Heating System: Supplementary Information: Calibrating the laser heating system. The laser control system with all its components is showcased in Figure S1.
We implement a pulse-width-modulation algorithm on the Teensy-4 microcontroller to interpolate the intensity value in the range [01][0-1] using the maximum operational current of 5A5\ A and the minimum packet communication latency of 50ms50\ ms. Since most of the control actions rely on shrinking the muscles to a target length, we chose the configuration of laser parameters that allow us to actuate the maximum number of muscles concurrently. The experiment details are explained in Constraints on the Number of Actuators in a Robot: Supplementary Information: Constraints on the Number of Actuators in a Robot. With this configuration, we can achieve a total muscle contraction length of up to 300mm300\ $m$. For a standard muscle of length 30mm30\ $m$, this corresponds to contracting approximately 30 muscles, each by up to 33%. This total shrink length and contraction ratio define the fabrication limitations of the skeletal designs imposed by the control system (see Equation 2). For this configuration, we further characterize muscle shrinkage by varying the number of repetitions of laser actuation on an LCE muscle of length 30mm30\ $m$. Figure 3 (D) showcases the shrinkage behavior of muscles when activated with the laser setup for maximum intensity while varying the number of repetitions.
@@
1:Input: 3D manifold meshes $\{\Omega^{1}_{0}(V^{1}_{0},F^{1}_{0})$, $\Omega^{2}_{0}(V^{2}_{0},F^{2}_{0})$$\Omega^{n}_{0}(V^{n}_{0},F^{n}_{0})\}$
2:Output: 3D approximated meshes $\{\Omega^{1\star}(V^{1\star},T^{\star}),\Omega^{2\star}(V^{2\star},T^{\star})...\Omega^{n\star}(V^{n\star},T^{\star})$}] and 2D mesh $\Gamma^{\star}(V^{flat\star},T^{\star})$, all sharing the same triangulation $T^{\star}$.
3:Energy Function: for the current meshes $\{\Omega^{i}(V^{i},T)\}$, $\Gamma(V^{flat},T)$,
$\mathcal{E}=\frac{1}{{N\choose 2}}\sum_{ij}E_{map}(\Omega^{i},\Omega^{j})\ +\ \frac{1}{N}\sum_{i}E_{mesh}(\Omega^{i})\ +\ \frac{1}{N}\sum_{i}E_{fab}(\Gamma,\Omega^{i}).\quad$ (1)
Constraints: $\forall i\in[1,n]$ and $\forall ej\in T$, and $\mathcal{L}$ be the length of edge $e_{j}$,
@@
$\forall_{e_{j}\in T}:\mathcal{L}\left(\Gamma(e_{j})\right)>30.$ (2)
4:Initialize a coarse triangulation $T$ (two triangles) and then initialize $\Omega^{i}(V^{i},T$) by selecting extreme boundary vertices of $\Omega^{i}_{0}$ as $V^{i}$. Initialize $\Gamma(V^{flat0},T)$ with $V^{flat0}$ as vertices on the boundary of unit circle.
5:Optimization: The common triangulation $T$ and the vertex positions $\{V^{i}\}$ and $V^{flat}$ are optimized by minimizing the energy function (Algorithm Details of Computational Design for Shape Morphing Structures) while adhering to the the fabrication constraints (Equation 2)
@@
The design tool aims to generate the skeletal graph for robots that can morph between different target deformations. It takes the desired target shapes as input and generates a flat skeletal graph, which after fabrication, can be actuated by the laser heating system to closely approximate the input target shapes. To obtain the skeletal graph, we approximate the target shapes with the same connectivity and ensure their edge lengths are within the shrinkable range of the flat rest graph. This allows our edge-shrinkage-based method to morph between them. We adapt the framework proposed by  (?) aimed at computing surface homeomorphisms by optimizing compatible triangulations for various objects. This triangulation when embedded in 2D yields a planar skeletal graph structure. We summarize our method in 1. The detailed modifications made to the formulation proposed by  (?) and the optimization steps are explained in Supplementary Information: Algorithmic Details of Computational Design for Shape Morphing Structures. We show the efficacy of our tool by building the morph-bot (see Figure 5), which morphs into target saddle and dome shapes with good visual accuracy. We further showcase futuristic applications that the computational design tool unlocks by fabricating finely tessellated flat-fabricated surfaces that morph into multiple target configurations (see Figure S2).
@@
The numerical simulation uses the robot’s skeletal graph as input. We model the bones of the skeletal graph using discrete elastic rods  (?). The LCE muscles are modeled as spring actuators in the simulation, with their rest length being a function of temperature. We adjusted the material parameters of the simulation model to ensure that the motion of the base building block matches its behavior in the real world. This is done by tuning the simulation’s material parameters as described in  (?). We mimic the muscle shrinkage as result of laser actuation, by gradually changing the rest length of the spring actuators from their initial length to the target length during the simulation. Finally, we use a quasi-static simulation framework with contact and gravity to obtain the final shape of the robots after morphing. This allows us to verify the efficacy of the shape-morphing robots before they are manufactured. Figure 5 (C) shows the simulation results of morph-bot .
@@
We introduce a design tool to fine-tune a given skeletal robot to achieve a desired functional task. The design space includes the length of active actuators and their control policy. We jointly optimize these two spaces to obtain a skeleton with a control scheme that performs the target function. If multiple functions are desired, we output one unique optimized skeleton and a control scheme for each function. We model the robots as rigid articulated robots and simulated them using  Mujoco  (?). We simulate three different target behaviors: walking on flat terrain, climbing on an inclined plane, and swimming in water. Since  Mujoco allows for fast simulation of rigid robots, we can optimize the nodal positions of the skeleton and control schemes for different target behavior using evolutionary search based algorithms. The implementation details of our optimization framework are described in Supplementary Information: Algorithmic Details for exploring the Design and Control of Terrain Navigating Robots. The optimized skeletons exhibit the highest fitness scores for their respective target behavior. We use this tool for the designs of frog-bots . Modeling the robots with rigid linkages introduces a significant sim-to-real gap, leading to discrepancy between simulated and fabricated motions. We use the skeletal design produced by our tool and manually tune the actuation profiles for the frog-bots to maximize their performance in the real world. Despite this, the design tool successfully optimizes the skeletal graph of frog-bots for their specific tasks, outperforming the non-optimized skeleton frog-bot . Figure 6 and supplementary video: S6, showcase the frog-bots designed using this tool.